### v2

# Further author information: (Send correspondence to Chelsea Adelman)

Abstract
Kernel phase interferometry (KPI) is a post-processing technique that treats a conventional telescope as an interferometer by accurately modeling a telescope pupil as an array of virtual subapertures. KPI provides angular resolution within the diffraction limit by eliminating instrumental phase errors to first order. It has been successfully demonstrated to boost angular resolution on both space- and ground-based observatories, and is especially useful for enhancing space telescopes, as their diameters are smaller than the largest ground-based facilities. Here we present the first demonstration of KPI on JWST/MIRI data at 7.7 microns, 10 microns, and 15 microns. We generate contrast curves for 16 white dwarfs from the MIRI Exoplanets Orbiting White dwarfs (MEOW) Survey, finding significantly deeper contrast at small angular separations compared to traditional imaging with JWST/MIRI, down to within /D. Additionally, we use our KPI setup to successfully recover four known companions orbiting white dwarfs and brown dwarfs. This analysis shows that at these wavelengths KPI can uniquely access the orbital parameter space where inward-migrating post-main-sequence giant exoplanets are now thought to exist. We discuss the prospects for applying KPI to a larger sample of white dwarfs observed with JWST, increasing the volume of directly imaged close-in post-main-sequence exoplanets.
keywords:
kernel phase interferometry, JWST/MIRI, exoplanets, white dwarf, brown dwarf, post main sequence evolution, planetary evolution1 INTRODUCTION
Kernel phase interferometry (KPI) is a post processing technique that has been successful in resolving known companions, and in detecting unknown companions unattainable with standard imaging, on both space-based [16, 6, 1] and ground-based observatories[17, 7, 5]. By modeling a conventional telescope as a densely-packed interferometer, KPI accesses angular separations that extend within the diffraction limit of the telescope [14, 20, 4]. It is best used for moderate contrast imaging, and requires high Strehl (0.85)[20]. KPI has also been successfully demonstrated on JWST/NIRISS [6], but has not yet been implemented on other JWST instruments. The combination of JWST’s stability, high Strehl, and long wavelength would make long-wavelength JWST KPI useful for detecting colder companions than those accessible from the ground or with HST KPI.
Here we present the first demonstration of KPI on JWST/MIRI. KPI’s moderate contrast makes it well suited for constraining the occurrence rates of post-main-sequence (post-MS) planets around white dwarfs (WDs), which has recently been demonstrated as an exciting JWST science application [21, 9, 10, 18, 15]. JWST’s sensitivity also makes KPI a viable technique for companion searches around faint brown dwarfs (BDs). As a proof of concept, we utilize archival data of white dwarfs and brown dwarfs for KPI analysis. We create achievable contrast curves for 16 of the 17 targets in the MIRI Exoplanets Orbiting White dwarfs (MEOW) survey (PID 4403). We also attempt to recover known companions orbiting two BDs, 014656.66+423410.0[3] and J171104.60+350036.8[11] (referred to as W0416 and W1711, respectively, hereafter), which were observed as part of a JWST cycle one program (PID 2124). Additionally, we attempt to recover the two first directly imaged exoplanets orbiting WDs, WD 1202-232 and WD 2105-82[15], which were discovered via another JWST cycle one program (PID 1911).
In this manuscript, we discuss KPI in greater detail in Section 2 and the interferometric model creation and parameters in Section 3. We then describe the data and the reduction techniques in detail in Section 5. We present the results in Section 6 and discuss their implications as well as future plans in Section 7. Lastly, we conclude in Section 8.
2 Kernel Phase Interferometry
Kernel phase interferometry treats a conventional telescope like an interferometer [14]. The amplitudes and phases of complex visibilities are calculated via the Fourier transform (FT) of the images. The filled aperture is modeled as a redundant interferometric array with a linear relationship between instrumental phases (which are assumed to be small; ) and Fourier phases. Instrumental errors can be eliminated to the first order by using a linear combination of the Fourier phases, called kernel phases (KPs).
The following equations describe the the kernel phase calculations. Equation 1 describes the linear model that connects the pupil plane phases and the Fourier phases.
Here is a vector of measured complex visibility phases, and is a square diagonal matrix consisting of redundancy values of the baselines. is a matrix representing the pairs of subapertures contributing to each spatial frequency in the FT, and represents the instrumental phases associated with the subapertures in the pupil plane. Multiplying both sides by gives
To eliminate the instrumental errors, the left null space of , called , is formulated. Left multiplying on both sides of Equation 2 eliminates the instrumental term and leaves the intrinsic phase information of the target. This is shown in Equation 3.
3 KPI Pupil Model Generation
Constructing the matrix requires a pupil model, which we generate using the Python package now known as the Space Telescope Point Spread Function (STPSF111https://github.com/spacetelescope/stpsf) formerly known as webbpsf. This package calculates simulated PSFs for JWST. We use the STPSF JWST/MIRI pupil model with XARA222https://github.com/fmartinache/xara/, a python package designed for KPI calculations. XARA can be used to create “grey” or “binary” models. Here, binary means transmission is either 0% or 100% for all subpertures throughout all parts of the pupil. Grey models account for diffraction that occurs around the spiders, secondary mirror, mirror segment gaps, and edges of the primary mirror, where transmission will be between 0% and 100%. Using XARA, we create a grey interferometric model and set the transmission cutoff (the transmission value where subapertures would be excluded) to be 0.70, previously used with KPI on JWST/NIRISS [6]. This model can be found in Figure 1.
We define the pupil model pitch, , or the spacing between each subaperture, as 0.35m. The pitch is chosen by visual inspection of the (u,v) points where sampling takes place in the observed Fourier-transformed images (see Section 2). It is chosen to result in sampling dense enough to resolve phase variations in the FT. Since the pitch determines the lowest sampled spatial frequency, when using KPI to constrain an object’s morphology, the effective field of view (FoV) is set by the pitch and wavelength, where the FoV . As described in Section 5, it is also good practice to use this effective field of view to set the minimum cropped image size. We use this pupil model to perform a KPI analysis of WDs and BDs observed in various filters (see Section 4), and discuss strategies for optimizing its parameters in Section 7.
4 Observations
The data in this work come from the publicly available Barbara A. Mikulski Archive for Space Telescopes (MAST) and they were collected using JWST’s Mid-Infrared Instrument (MIRI). We use three different datasets from three different JWST programs (PIDs: 4403, 1911, and 2124) to perform a MIRI KPI proof of concept. We analyze 18 WDs and two BDs. The observations for each dataset are described below.
The MEOW Survey (PID 4403): We reduce data taken as part of the MEOW Survey with the goal of generating contrast curves. The data were collected from September 9, 2023 to May 4, 2024 and comprise 17 WDs observed with JWST/MIRI. All targets were observed in F770W, F1800W, and F2100W, corresponding to 7.7 m, 18 m, and 21 m. The exposure times for each target in the F770W, F1800W, and F2100W filters are 55.5s, 277.504s, and 710.412s, respectively. One target, G 107-70, is in a tight binary, so we chose to exclude it from our analysis. We focus on the other 16 targets, specifically those observed in the F770W filter.
Brown Dwarfs: W0146 and W1711 (PID 2124): We reduce the W0146 and W1711 data with the goal of recovering known, close-in companions. The data for W1711 and W0146 were collected on May 31, 2023 and September 21, 2022, respectively. W0146 has a known companion orbiting at mas, E of N with a contrast of mag, and W1711 has a known companion orbiting at mas, E of N with a contrast of mag [2]. Both targets are observed with F1000W, F1280W, and F1800W filters corresponding to 10 m, 12.8 m, and 18 m, respectively. Exposure times for F1000W, F1280W, and F1800W images are 38.85s, 44.4s, and 55.5s, respectively for W0146. Exposure times for F1000W, F1280W, and F1800W images are 27.75s, 38.85s, and 55.5s, respectively for W1711. We focus specifically on the data collected with the F1000W filter.
White Dwarfs: WD 1202-232 and WD 2105-82 (PID 1911): We analyzzed observation of WD 1202-232 and WD 2105-82 to recover known wider separation companions. The data were collected for WD 1202-232 and WD 2105-82 on February 9, 2023 and April 21, 2023, respectively. WD 1202-232 has a known companion orbiting at ′′, E of N with a contrast of mag, and WD 2105-82 has a known companion orbiting at ′′, E of N with a contrast of mag. Both targets were observed in F560W, F770W, F1500W, and F2100W, corresponding to 5.6 m, 7.7 m, 15 m, and 21 m respectively. The exposure times for F560W, F770W, F1500W, and F2100W data are 255.304s, 277.504s, 8413.92s, and 1309.816s respectively for WD 1202-232. The exposure times for F560W, F770W, F1500W, and F2100W images are 233.104s, 233.104s, 12088.08s, and 6016.288s respectively for WD 2105-82. We specifically focus on the F1500W data for these two targets.
5 Data Reduction
We first use aperture photometry to perform background subtraction for each target. To do this we use a Python package called photutils. The median of an annulus around each target is calculated to measure local background noise, then subtracted from the image. Each image is then cropped. As described in Section 3, given the pupil sampling pitch, the KPIs would be sensitive to spatial scales as large as 4.5′′x4.5′′, 5.8′′x5.8′′, and 8.8′′x8.8′′ at 7.7 m, 10 m, and 15 m, respectively. Each MEOW target (observed at 7.7 m) is cropped to 40x40 pixels, corresponding to a FoV of 4.4′′x4.4′′. The BD data (obtained at 10 m) are cropped to 50x50 pixels, corresponding to a FoV of 5.5′′x5.5′′. Both of the images with WDs (observed at 15 m) are cropped more aggressively because of faint background sources near the science target. WD 2105-82 is cropped to 64x64 pixels, corresponding to a FoV of 7′′x7′′, and WD 1202-232 is cropped to 30x30 pixels, corresponding to a FoV of 3.3′′x3.3′′. The downside to cropping aggressively is losing the sampling of some of the shorter baselines, or lower spatial frequencies, but because the separations of these companion are small compared to the lowest spatial frequencies, this did not inhibit recovering the signal.
The image is then corrected for bad pixels by comparing each pixel to a 10 x 10 surrounding box. A bad pixel is flagged and replaced with the median of its unflagged neighbors if its value exceeds the box’s average by more than . We then apply a super-Gaussian window to the image to eliminate noisy pixels far from the PSF center. Without this suppression, these noisy pixels can create correlations in the FT. Most of the targets have a super-Gaussian of order four and HWHM defined as converted to pixels (, , and pixels for 7.7 m, 10 m, and 15 m, respectively). We are aggressive with the windowing of W0146, with a HWHM of 20 pixels, as noisy pixels far from the PSF were mimicking a signal. Then using XARA and the pupil model, we take the FT of the images and sample the FT at (u,v) coordinates to extract KPs, as shown in Figure 2. We assign conservative error bars by measuring the standard deviation of all of the final KPs and assigning it as a uniform error to every KP.
The data come from observing programs that were not ideal for performing good calibration because each target is observed once in each filter, and no reference stars were observed. So, for this proof of concept, all calculated KPs are raw and uncalibrated. Typically for calibration we use either angular differential imaging [12][8] if we have images observed at two roll angles, or reference star differential imaging (RDI)[7] if unresolved objects were observed close in time to the science target. Interferometric calibration using RDI on JWST requires calibrators to be observed as close as possible on the detector and with peak counts as close as possible [19]. One of the next steps is to carefully calibrate the KPs for optimization of companion detectability, discussed further in Section 7.
We next use XARA functions to create a set of KPs for binary models of various combinations of separations, position angles, and contrasts, creating a three dimensional grid of model KPs for all parameter combinations. The best fit binary model is found by minimizing the calculated from the model KPs and the data. If no significant signal is detected, we then use the same grid of companion fits to generate contrast curves. For these, values are averaged across the position angles to create a grid of as a function of separation and contrast. We use intervals of values between the null model, with a separation of 0 mas, and the companion models to determine the detectability confidence levels. We create the final contrast curves by choosing contrasts with of 25, corresponding to being distinguishable from the null model at [20].
6 Results
Here we present the contrast curves and companion fitting results. We create contrast curves for 16 of the 17 targets in the MEOW dataset, shown in Figure 3. We achieve contrasts of mag for most MEOW targets within ( mas or AU at the average MEOW target distance of pc away). Standard imaging can achieve contrast of roughly 6 mag at mas for 5.8 m, and deeper contrasts at mas [18] (assuming similar performance, this translates to mas scaled for a wavelength 7.7 m). These results demonstrate that MIRI KPI provides a boost in angular resolution, decreasing the inner working angle by a factor of .
| ID | SepDI | SepKPI | PADI | PAKPI | ContrastDI | ContrastKPI |
| [mas] | [mas] | [deg] | [deg] | [mag] | [mag] | |
| J014656.66+423410.0 | ||||||
| J171104.60+350036.8 | ||||||
| WD1202-232 | ||||||
| WD2105-82 |
All KPI recovered parameters and DI reported parameters for the four companion recoveries can be found in Table 1. Figure 4 shows the images and DI/KPI companions positions for the two BDs (W0146 and W1711) and two WDs (WD1202-232 and WD2105-82). As shown in Table 1, all DI and KPI companion parameters for W1711 have error bars that overlapped. For WD1202-232 and WD2105-82, recovered separations and contrasts have overlapping error bars, while the position angles differ by slightly more than . The differences between the KPI and DI best fit parameters for W0146 are significantly larger than the parameter error bars, at 40.23 mas, 37.1 degrees, and 0.991 mag in separation, position angle, and contrast, respectively. These discrepancies could be caused by systematic errors present in the uncalibrated KPs, which would bias the companion fits and also lead to underestimated parameter error bars (since we assume Gaussian KP errors). We generally expect that kernel phase calibration may cause the KPI outputs to converge to similar values as those reported for the DI analysis of the BD companions [2] and WD companions [15].
The KPs associated with the best fit binary parameters are compared with the KP data in Figure 5. If a companion were present with high SNR, the data would lie along the black dashed line, indicating a one-to-one correlation. We see that W1711 and WD 1202-232 have a strong correlation with the best fit binary model, but W0146 and WD 2105-82 are not as strongly correlated. This may be due to higher order systematic noise, which can be improved with kernel phase calibration (see Section 7).
7 Discussion
We achieve contrasts of mag for most of the MEOW targets at mas. The WD distances in this sample range from pc, with an average distance of 12au. The angular resolution of 200 mas (which is within for JWST at m; mas) at these distances thus translates to au. Previous analysis of WD 2149+021 using standard imaging on JWST/MIRI led to a mass sensitivity of M at the inner most working angle of 654 mas (corresponding to a contrast of at 6.5 mag at 5.8 m) [18]. Assuming similar performance as presented here, observing this system with KPI would result in a slightly lower mass sensitivity (reaching 6 mag of contrast compared to 6.5 mag), at an inner working angle of mas, within . KPI’s mass sensitivity at these separations can help probe the region where inward migrating post-MS giants are thought to reside [22, 21, 9, 10]. We expect this sensitivity to improve when robust calibration strategies are applied.
We successfully recover known companions in WD and BD systems with uncalibrated KPs, with separations ranging from mas to ′′ corresponding to . The KPI parameters for three of the companions are in agreement with the DI parameters previously reported [2, 15]. Likely due to the use of uncalibrated data, the parameters for W0146 (while significant compared to the null model) were discrepant with the DI analysis. Calibration has been shown to decrease the scatter of KPs by removing residual systematic errors[20, 6], which may be required for robust recovery of close-in () companions like the one orbiting W0146. Despite these systematic offsets for the closest separations, these companion recoveries demonstrate that even uncalibrated KPI can enable JWST super-resolution. They also highlight the prospects for future calibrated KPI MIRI investigations.
Improvements to the MIRI pupil model and exploration of calibration options would benefit MIRI KPI and will be the subject of future work. Adjusting the pupil sampling pitch and transmission cutoff would optimize the pupil model. Decreasing the pupil sampling pitch () would increase the density of the Fourier phase sampling. We plan to explore various pitch settings to find the minimum viable value, determined by the where adjacent (u,v) sampling points no longer add new information. For transmission cutoff, previous KPI studies on JWST/NIRISS used a value of 70% [6], and future work will determine the best cutoff for MIRI. The methodology in Martinache et al. (2020)[13] can be used to optimize both pitch and transmission cutoff. The expected diagonal redundancy matrix can be compared to the amplitude of the complex visibilities extracted from the data, and they should match for an optimized pupil.
Future work will also determine robust kernel phase observational strategies and calibration methods. Ideally, RDI KPI datasets would include PSF calibrator observations obtained close in time to the science target observations, on the same region of the detector. While testing the limits of KPI calibration may require a dedicated MIRI program, archival datasets with many objects observed close in time could be used to test calibration quality as a function of relative detector position, peak counts, and wavefront drift. Similar comparisons could be performed using multiple PSFs in the same field as the science target. Careful dedicated and archival calibration tests will inform future KPI proposal planning and re-analysis of archival data that was originally processed with traditional methods.
8 Conclusion
We present the first demonstration of KPI on JWST/MIRI. We use 16 white dwarfs in the MEOW survey to produce achievable contrast curves, where we are sensitive to companion contrasts of magnitudes for most of the targets at mas. This angular separation is within on JWST at m, and corresponds to an orbital separation of au, given the average MEOW target distance. Additionally, we are able to recover known companion signals in WD systems and BD systems, with separations of mas and contrasts of mag.
This KPI demonstration uses raw KPs, as well as a pupil sampling pitch chosen by visually inspecting the (u,v) coverage and a transmission cutoff chosen from previous KPI work done on a different JWST instrument. We expect the KPI sensitivity to improve with KP calibration, as well as with optimization of the interferometric model pitch and transmission cutoff. Even without calibration, the achievable contrasts presented here reach giant planet masses for WD and BD hosts. This demonstrates the utility of raw KPI for substellar companion searches, motivating its use for re-processing archival data originally processed with traditional methods. Improvements enabled by future work will access even fainter and thus colder / lower-mass companions, making JWST/MIRI KPI a promising technique for expanding the directly-imaged exoplanet parameter space.
Acknowledgements.
We would like to thank Cal Bridge for providing financial support for C.A. graduate student research. M.D.F. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303911. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs 4403, 2124, and 1911. All of the data presented in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via DOI 10.17909/zcr4-9k86. STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5–26555. Support to MAST for these data is provided by the NASA Office of Space Science via grant NAG5–7584 and by other grants and contracts.References
- [1] (2024-08) Planet search around seven white dwarfs in the Hyades cluster using kernel phase interferometry. In Optical and Infrared Interferometry and Imaging IX, J. Kammerer, S. Sallum, and J. Sanchez-Bermudez (Eds.), Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 13095, pp. 130952V. External Links: Document Cited by: §1.
- [2] (2025-08) Discovery of the Second Y+Y Dwarf Binary System: CWISEP J193518.59-154620.3. arXiv e-prints, pp. arXiv:2508.17176. External Links: Document, 2508.17176 Cited by: §4, Table 1, §6, §7.
- [3] (2015-04) Discovery of a Low-luminosity, Tight Substellar Binary at the T/Y Transition. The Astrophysical Journal 803 (2), pp. 102. External Links: Document, 1502.04707 Cited by: §1.
- [4] (2013-08) Phase errors in diffraction-limited imaging: contrast limits for sparse aperture masking. Monthly Notices of the Royal Astronomical Society 433 (2), pp. 1718–1728. External Links: Document, 1301.6205 Cited by: §1.
- [5] (2014-01) Orbital Motion and Multi-Wavelength Monitoring of LkCa15 b. In Exploring the Formation and Evolution of Planetary Systems, M. Booth, B. C. Matthews, and J. R. Graham (Eds.), IAU Symposium, Vol. 299, pp. 199–203. External Links: Document Cited by: §1.
- [6] (2023-01) The Near Infrared Imager and Slitless Spectrograph for JWST. V. Kernel Phase Imaging and Data Analysis. Publications of the Astronomical Society of the Pacific 135 (1043), pp. 014502. External Links: Document, 2210.17528 Cited by: §1, §3, §7, §7.
- [7] (2019-06) Kernel phase imaging with VLT/NACO: high-contrast detection of new candidate low-mass stellar companions at the diffraction limit. Monthly Notices of the Royal Astronomical Society 486 (1), pp. 639–654. External Links: Document, 1903.11252 Cited by: §1, §5.
- [8] (2020-04) Angular differential kernel phases. Astronomy and Astrophysics 636, pp. A21. External Links: Document, 2003.04020 Cited by: §5.
- [9] (2025-05) Thermal Emission and Confirmation of the Frigid White Dwarf Exoplanet WD 1856+534 b. The Astrophysical Journal Letters 984 (1), pp. L28. External Links: Document, 2504.16982 Cited by: §1, §7.
- [10] (2024-09) The MIRI Exoplanets Orbiting White dwarfs (MEOW) Survey: Mid-infrared Excess Reveals a Giant Planet Candidate around a Nearby White Dwarf. The Astrophysical Journal Letters 973 (1), pp. L11. External Links: Document, 2408.16813 Cited by: §1, §7.
- [11] (2012-10) Two Extraordinary Substellar Binaries at the T/Y Transition and the Y-band Fluxes of the Coolest Brown Dwarfs. The Astrophysical Journal 758 (1), pp. 57. External Links: Document, 1206.4044 Cited by: §1.
- [12] (2006-04) Angular Differential Imaging: A Powerful High-Contrast Imaging Technique. The Astrophysical Journal 641 (1), pp. 556–564. External Links: Document, astro-ph/0512335 Cited by: §5.
- [13] (2020-04) Kernel-phase analysis: Aperture modeling prescriptions that minimize calibration errors. Astronomy and Astrophysics 636, pp. A72. External Links: Document, 2003.02032 Cited by: §7.
- [14] (2010-11) Kernel Phase in Fizeau Interferometry. The Astrophysical Journal 724 (1), pp. 464–469. External Links: Document, 1009.3933 Cited by: §1, §2.
- [15] (2024-01) JWST Directly Images Giant Planet Candidates Around Two Metal-Polluted White Dwarf Stars. arXiv e-prints, pp. arXiv:2401.13153. External Links: Document, 2401.13153 Cited by: §1, Table 1, §6, §7.
- [16] (2013-04) Dancing in the Dark: New Brown Dwarf Binaries from Kernel Phase Interferometry. The Astrophysical Journal 767 (2), pp. 110. External Links: Document, 1302.6682 Cited by: §1.
- [17] (2016-01) The Palomar kernel-phase experiment: testing kernel phase interferometry for ground-based astronomical observations. Monthly Notices of the Royal Astronomical Society 455 (2), pp. 1647–1653. External Links: Document, 1510.06406 Cited by: §1.
- [18] (2024-06) A MIRI Search for Planets and Dust around WD 2149+021. The Astronomical Journal 167 (6), pp. 257. External Links: Document, 2311.14165 Cited by: §1, §6, §7.
- [19] (2024-03) The JWST Early Release Science Program for Direct Observations of Exoplanetary Systems. IV. NIRISS Aperture Masking Interferometry Performance and Lessons Learned. The Astrophysical Journal Letters 963 (1), pp. L2. External Links: Document, 2310.11499 Cited by: §5.
- [20] (2019-01) Comparing nonredundant masking and filled-aperture kernel phase for exoplanet detection and characterization. Journal of Astronomical Telescopes, Instruments, and Systems 5, pp. 018001. External Links: Document, 1901.01266 Cited by: §1, §5, §7.
- [21] (2020-09) A giant planet candidate transiting a white dwarf. Nature 585 (7825), pp. 363–367. External Links: Document, 2009.07282 Cited by: §1, §7.
- [22] (2015-02) Detectable close-in planets around white dwarfs through late unpacking. Monthly Notices of the Royal Astronomical Society 447 (2), pp. 1049–1058. External Links: Document, 1411.6012 Cited by: §7.